OpenStudy (anonymous):

find the second derivative using implicit differentiation: 9x^2 + y^2 = 9

OpenStudy (anonymous):

(9x^2)dy/d + (y^2)dy/d = (9)dy/d 18x + 2y(y') = 0 y' = -18x/2y

OpenStudy (anonymous):

or y' = -9/y

OpenStudy (anonymous):

9x^2 + y^2 = 9, differentiating to get 18x+2y y '=0,differentiating again to get 18+2[y y ''+y 'y ']=0 [y y ''+y 'y ']=-18/2=-9 [y y ''+(y ')^2]=-9 y y ''=-(y ')^2-9 y ''=(-(y ')^2-9)/y ans....